### v1

# ShareChat: A Dataset of Chatbot Conversations in the Wild

Abstract
While Large Language Models (LLMs) have evolved into distinct platforms with unique interface designs and capabilities, existing public datasets treat models as generic text generators, stripping away the interface context that actively shapes user interaction. To address this limitation, we present ShareChat, a large-scale, cross-platform corpus comprising 142,808 conversations and over 660,000 turns collected from publicly shared URLs across five major platforms: ChatGPT, Claude, Gemini, Perplexity, and Grok. ShareChat distinguishes itself by preserving native platform affordances often lost in standard logs, including reasoning traces, source links, and code artifacts, while spanning 101 languages over the period from April 2023 to October 2025. Furthermore, ShareChat offers substantially longer context windows and greater interaction depth than prior datasets. We demonstrate the dataset’s multifaceted utility through three demonstrative use cases: (1) analyzing conversation completeness to measure user intent satisfaction; (2) evaluating source citation behaviors in content generation; and (3) conducting temporal analysis to track evolving usage patterns. This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild
[…]
ShareChat: A Dataset of Chatbot Conversations in the Wild
Yueru Yan Tuc Nguyen Bo Su Melissa Lieffers Thai Le Indiana University Bloomington, USA {yueryan,tucnguye,subo,mealieff,tle}@iu.edu
1 Introduction
Conversational large language model (LLM)-based chatbot services have evolved rapidly in the past three years. The first widely adopted general purpose LLM chatbot, ChatGPT, was launched in November 2022 and has reached more than 700 million weekly active users by mid 2025 OpenAI Research (2025). Following this success, many companies quickly released their own chatbot services, which also attracted substantial public attention. For example, Anthropic introduced the Claude families in 2023 Anthropic (2023, 2024, 2025), Google deployed the Gemini family of models Anil et al. (2023), xAI launched Grok as a chatbot integrated with the social media platform X xAI (2023), Perplexity emerged as an answer engine that combines conversational interaction with web search Perplexity AI (2024), and DeepSeek released its R1 reasoning model and corresponding consumer chatbot app in January 2025 DeepSeek AI (2025). Although all of these services are built on text-based LLMs, they differ in interface design, supported features, and safety policies, which in turn shape how users interact with them.
For example, Grok can surface live posts from X when providing answers, Claude models are optimized and evaluated for coding, math, and analysis tasks Anthropic (2023, 2024), and Perplexity consistently presents responses with explicit source citations Perplexity AI (2024). These platform level design choices are not static. Commercial systems typically undergo continuous reinforcement learning and fine tuning on conversation logs and human feedback Ouyang et al. (2022); Bai et al. (2022); Chen et al. (2024), which tends to reinforce platform specific strengths and norms over time. As a result, it is important to understand interaction dynamics and user experience across platforms, instead of treating LLM capabilities as a property of models in isolation.
However, most widely used LLM benchmarks focus on curated single turn tasks, such as multiple choice knowledge tests or coding and math word problems Hendrycks et al. (2021); Cobbe et al. (2021). In practice, the use of conversational LLMs is much more diverse and complex: users ask follow up questions, change goals, or co construct content over many turns. Recent work has shown that even state of the art models often get “lost” in multi turn conversations, with reliability degrading as instructions are refined across turns Laban et al. (2025a). Studying how people actually interact with LLMs in the wild therefore requires large scale, real conversation datasets, rather than only synthetic prompts or short, benchmark style dialogues.
Several real world chatbot LLM conversation datasets have been released in the past two years Zhao et al. (2024); Zheng et al. (2024); Köpf et al. (2023). For instance, WildChat Zhao et al. (2024) compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs. LMSYS-Chat-1M Zheng et al. (2024) contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena). OpenAssistant Conversations Köpf et al. (2023) provides human generated assistant style dialogues and detailed preference annotations to support alignment research. These datasets are important resources, but they have several limitations:
Single-Interface Mediation: They are all mediated via a central interface even when multiple backend models are available. Consequently, they do not fully capture how different commercial platforms, with distinct interaction designs and product positioning, influence the content that users share and the way they engage with LLMs.
Text-Only Representation: In addition, existing real conversation datasets primarily focus on plain text message content. Yet real interactions with contemporary LLM systems can include more structure: users may enable a thinking mode that exposes intermediate reasoning traces, and some platforms attach source links or other metadata to model outputs, such as the cited web pages surfaced by Perplexity or embedded X posts shown by Grok. A richer multi platform dataset should preserve these elements when they are available, so that researchers can better understand how users interpret model reasoning, how they rely on external evidence, and how platform affordances shape behavior across turns.
Limited Interaction Depth: Moreover, although some datasets explicitly include multi turn conversations Zhao et al. (2024), the average depth of a conversation is still relatively short. Given that context length is central to LLM training and performance, longer real user chatbot conversations are valuable both for understanding how user intents and prompts evolve and for investigating when and how models succeed or fail at using extended context Laban et al. (2025a).
Observer Bias: Finally, data collection in these studies typically relies on pre-interaction consent, where users agree to share their data before or during the interaction in exchange for service access. This introduces the Observer Bias (or Hawthorne Effect) McCambridge et al. (2014), as users who are aware they are being monitored for research purposes may consciously or unconsciously alter their behavior, which often exhibits higher social desirability compared to authentic, unobserved private usage. In contrast, ShareChat relies on post-hoc sharing, where the decision to publish is made only after the natural interaction has concluded, thereby preserving the integrity of the original behavior during the conversation itself.
In this study, we address these gaps by constructing a large scale, multi platform corpus of authentic user-LLM conversations. We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: Claude, Perplexity, Grok, ChatGPT, and Gemini. All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than twenty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages. The dataset therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers. Each conversation record contains the complete sequence of user and assistant turns, along with platform specific metadata such as, when available, timestamps, model thinking traces, or source links attached by the interface. The data collection is done under IRB approval (#28569).
To demonstrate the specific utility of ShareChat, besides describing the basic statistics of the corpus, we present three illustrative case studies, each designed to leverage ShareChat’s unique features to overcome the limitation of prior resources. First, we perform conversation completeness analysis, where we label whether specific user intentions are fully addressed. This utilizes ShareChat’s high interaction depth to capture the resolution of complex goals often obscured in shorter, shallow exchanges. Second, we carry out source domain analysis that connects assistant responses to external resources surfaced by platforms like Perplexity and Grok. This leverages our preservation of non-textual metadata to evaluate how models ground their claims, which is a dimension absent in text-only datasets. Third, we conduct temporal analysis to examine response time and engagement patterns. This exploits our native turn-level timestamps to reveal authentic real-world interaction rhythms, free from the temporal distortions often introduced by observer bias in monitored settings. Together, these analyses highlight the value of a multi-platform, multi-turn, and multi-lingual conversation corpus for understanding real user experiences with LLMs.
Our contributions are twofold. First, we construct and analyze what is, to our knowledge, the first multi platform, multi turn, and multi lingual conversation dataset of user LLM interactions, with substantially longer conversations on average than prior corpora and lower levels of toxic content. Second, we provide three concrete use cases: conversation completeness analysis, temporal analysis, and source link analysis, which together demonstrate how the dataset can be used to study user interaction behavior in depth.
2 Data Collection
For the five differnet platfomrs, we developed systematic, platform specific data extraction pipelines to collect conversation content, metadata, and platform specific features from publicly accessible shared conversation URLs found on the Internet. Although we found that such URLs are often shared on social media platforms such as X, Reddit p(formerly Twitter), Reddit, and Discord, exhaustively monitoring these individual ecosystems presents significant scalability challenges. To tackle this, we utilized Internet archival services, specifically the Internet Archive (Wayback Machine), to search for and extract URLs matching specific patterns (e.g., “chatgpt.com/share/*” for ChatGPT), This archival-based discovery method is a well-established strategy for aggregating public digital resources, previously employed to compile datasets for web-scale text Raffel et al. (2020), historical news archives Hamborg et al. (2017), and corporate disclosure analysis Boulland et al. (2024). Then, for each platform, data collection covered a distinct time window, as summarized in Table 1.
| Platform | Share URL Format | Collection Period |
| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
| Claude | claude.ai/share/* | Not available |
For every shared conversation page, we used automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation. The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform specific metadata. When certain elements were only visible after user interaction, we scripted the corresponding actions. For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
Filtering of Identifiable Information.
We prioritized user privacy by anonymizing personally identifiable information (PII). We used Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy222https://spacy.io/ for Named Entity Recognition, and custom rules to identify and remove PII across various data types—such as names, phone numbers, emails, credit cards, and URLs—in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
| Dataset | #Convs | #Users | #Turns | Avg. Turns | #User Tok | #Chatbot Tok | #Langs |
| Existing Public Datasets | |||||||
| Alpaca | 52,002 | – | 52,002 | 1.00 | 1 | ||
| Open Assistant | 46,283 | 13,500 | 108,341 | 2.34 | 11 | ||
| Dolly | 15,011 | – | 15,011 | 1.00 | 1 | ||
| ShareGPT | 94,145 | – | 330,239 | 3.51 | 41 | ||
| LMSYS-Chat-1M | 1,000,000 | 210,479 | 2,020,000 | 2.02 | 65 | ||
| Our Multi-Platform Dataset | |||||||
| Multi-Platform (Total) | 142,808 | – | 660,293 | 4.62 | 101 | ||
| Per-Platform Breakdown | |||||||
| ChatGPT | 102,740 | – | 542,148 | 5.28 | 101 | ||
| Perplexity | 17,305 | 4,763 | 24,378 | 1.41 | 45 | ||
| Grok | 14,415 | – | 53,094 | 3.69 | 60 | ||
| Gemini | 7,402 | – | 36,422 | 4.92 | 47 | ||
| Claude | 946 | – | 4,251 | 4.49 | 19 | ||
3 Preliminary Analysis
The final dataset contains conversations from five platforms, featuring a wide range of languages, diverse user prompts and longer conversation with less toxic content compared to other datasets. In this section, we describe the basic statistics of the dataset and platform-specific attributes. We also present the results of our toxicity analysis and topic distribution.
3.1 Basic Statistics
In total, ShareChat has 142,808 conversations with more than 660,293 interaction turns. Table 2 presents comprehensive statistics comparing our multi-platform dataset with existing conversation corpora on number of turns, token length and languages used. The language detection was performed at message-level using lingua-py and then combined into conversation-level statistics in Table 2. Token counts were computed using the Llama-2 tokenizer for consistent cross-platform comparison Zhao et al. (2024).
In summary, ShareChat significantly exceeds existing corpora in conversation depth, averaging 4.62 turns compared to the 2.02 turns observed in LMSYS-Chat-1M. Similarly, it demonstrates greater linguistic diversity by spanning 101 distinct languages, whereas the most diverse competitor covers only 65. The model responses are also substantially longer, with a mean of 1115.30 tokens compared to 215.71 for LMSYS-Chat-1M, suggesting the presence of richer and more substantive interactions. ChatGPT comprises the largest portion of the corpus with 102,740 conversations, offering the broadest language coverage.
The key distinguishing characteristic of ShareChat is the substantially longer and more heterogeneous distribution of conversational turns compared to existing public corpora as shown in Figure 1. Whereas prominent benchmarks such as Alpaca, Dolly, LMSYS-Chat-1M, and WildChat exhibit median turn lengths of 1.0, our aggregated dataset achieves a median of 2.0 turns, indicating that conversations in our corpus are fundamentally more multi-turn in nature. This pattern is consistently maintained across individual platforms with results shown in Appendix A.1.
For language detection, we first identified the language of each message using the lingua library, and we include these per message labels in the final dataset. We then assigned a dominant language to each conversation as the most frequently detected language across the user turns. As shown in Figure 2, English accounts for the clear majority of conversations across all platforms, with Japanese as the second most frequent language and all other languages each contributing a much smaller share. Compared to previous datasets, ShareChat has a more balanced language distribution.
As ShareChat features significantly longer conversation turns and higher token counts per message, it addresses a critical gap in current model development pipelines, which often rely heavily on single-turn instruction tuning. By capturing extended interactions, the data supports training for long-context dependencies and exploring complex user behaviors that only emerge during deep, multi-turn exchanges.
3.2 Platform-Specific Interaction and Metadata Attributes
Different platforms reflect distinct design philosophies. ChatGPT and Gemini offer general-purpose conversational assistance. Perplexity focuses on search-oriented information retrieval. Grok targets social media-integrated question answering. Claude specializes in technical and reasoning support. These structural differences align with intended use cases and shape the types of user interactions preserved in ShareChat. Table 3 outlines the distinct data attributes captured across the five platforms, which vary significantly based on their functional priorities.
While ChatGPT and Gemini follow standard conversational formats, other platforms expose specialized content structures. Claude is unique in preserving “thinking blocks” and versioned code artifacts, enabling the analysis of intermediate reasoning and iterative coding processes. In contrast, Perplexity and Grok structure responses around retrieval. Perplexity organizes content into answers, sources, and images with inline citations, while Grok features a dedicated pane for web and social media sources (X posts). These structural differences allow for distinct lines of inquiry: technical workflow analysis for Claude versus citation accuracy and source integration studies for Perplexity and Grok.
The granularity of metadata also dictates the scope of possible temporal analysis. ChatGPT and Grok are the only platforms providing turn-level timestamps, which are essential for modeling conversation rhythm and response latency. ChatGPT further enriches this with model version identifiers (e.g., gpt-4). Conversely, Perplexity lacks turn-level timing but offers unique engagement metrics, including view counts, share counts, and last-updated timestamps, effectively treating conversations as evolving public artifacts rather than static logs.
| Feature/Field |
Claude |
Perplexity |
Grok |
ChatGPT |
Gemini |
| Basic Turn Structure | ✓ | ✓ | ✓ | ✓ | ✓ |
| Metadata Capture | ✓ | ✓ | ✓ | ✓ | ✓ |
| HTML/MHTML | ✓ | ✓ | ✓ | - | ✓ |
| Code Artifacts | ✓ | - | - | - | - |
| Thinking Blocks | ✓ | - | ✓ | - | - |
| Analysis Blocks | ✓ | - | - | - | - |
| Source Citations | - | ✓ | ✓ | - | - |
| Image Metadata | - | ✓ | - | - | - |
| Timestamps / turn | - | - | ✓ | ✓ | - |
| User Identifiers | ✓ | ✓ | - | ✓ | ✓ |
| View Counts | - | ✓ | - | - | - |
3.3 Toxicity Analysis
To assess toxicity in ShareChat, we employed both DetoxifyHanu and Unitary team (2020) and OpenAI Moderation API333https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat analysis Zhao et al. (2024). Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering. Given the language coverage limitations of Detoxify, we retained only conversational turns in languages explicitly supported by the model, resulting in a filtered dataset of 949,487 conversations (72.9% of the original corpus) across all five platforms, with more details shown in Appendix A.2. For Detoxify, we applied the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
The results for turn-level statistics are shown in Figure 3. Our analysis revealed substantial differences between the two detection methods, consistent with limited agreement observed in prior work where WildChat Zhao et al. (2024) reported that Detoxify flagged 8.12% of user turns compared to the 6.05% flagged by OpenAI Moderation. At the turn level, we observed distinct patterns where Detoxify consistently detected higher toxicity rates for user messages, ranging from 2.8% to 5.6%, compared to the 1.3% to 4.0% range identified by OpenAI Moderation across all platforms. Notably, Claude showed the highest user toxicity under Detoxify detection at 5.6%. There appears to be a positive correlation between user and LLM toxicity rates within platforms. For example, Claude exhibited the highest user toxicity with 5.6% via Detoxify and 2.7% via OpenAI, and similarly showed elevated LLM toxicity at 6.4% and 3.1% respectively. Conversely, Perplexity demonstrated the lowest rates for both users, at 2.8% via Detoxify and 1.3% via OpenAI, and LLMs, at 0.5% via Detoxify and 1.8% via OpenAI.
Notably, ShareChat contains substantially less toxic content compared to existing benchmarks. Our overall user toxicity rates of 4.1% via Detoxify and 2.9% via OpenAI are considerably lower than both WildChat at 6.05% and LMSYS-Chat-1M at 3.08% using the OpenAI method Zhao et al. (2024). Similarly, our LLM toxicity rates of 1.6% via Detoxify and 3.2% via OpenAI are lower than the 5.18% reported by WildChat and comparable to the 4.12% in LMSYS-Chat-1M.
3.4 Topic Analysis
We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 444https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup. To enable large-scale processing, we run inference with 4-bit quantization, batch size 32. Model outputs are normalized through a post-processing pipeline to handle formatting variants and stray meta-responses, ensuring reliable mapping to our fixed taxonomy. For downstream analyses, the 24 fine-grained categories are further consolidated into seven higher-level groups to support both granular and coarse-grained comparisons across platforms. The 7 high-level categories of user requests are plot in the Figure. 4.
Overall, the averaged distribution reveals that information seeking remains the core driver of LLM interactions, with creative, technical, and advisory uses forming important but secondary dimensions of user behavior. Specifically, we observe a strong central tendency in how users engage with large language models. Seeking Information overwhelmingly dominates usage, comprising 39.6% of all requests, underscoring that fact-oriented question-answering remains the primary interaction paradigm. The second-largest category, Other/Unknown (19.0%), reflects a substantial proportion of prompts that are either ambiguous, exploratory, or insufficiently specified for reliable categorization. Technical Help (12.4%) and Writing (10.1%) constitute the next major cluster of user intent, highlighting the continued reliance on LLMs for procedural support and text generation tasks. Practical Guidance (8.4%) and Self-Expression (8.1%) appear at comparable frequencies, indicating moderate but consistent engagement with step-by-step advice and personal or expressive communication. Finally, Multimedia accounts for only 2.4% of requests, likely reflecting limited multimodal capabilities at the time of data collection. Detailed prompts, category definitions, and classification statistics are provided in the Appendix A.3.
4 Use Cases
4.1 Conversation Completeness Analysis
To assess how comprehensively large language models fulfill user needs in multi-turn help-seeking dialogues, we developed an automated completeness analysis pipeline that evaluates whether each user intention across a conversation is adequately addressed by the assistant. Our analysis reveals distinct performance across platforms, with ChatGPT and Claude demonstrating the highest consistency in fully satisfying user intentions, while Gemini, Grok and Perplexity exhibit significantly wider variance in conversational quality across multi-turn interactions.
The pipeline implements a three-stage workflow adapted from the DeepEval555https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework. First, we extract user intentions from each conversation by prompting Qwen3-8B Yang et al. (2025) with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation. Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress). Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0. This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes. We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions. The prompts used and experiment details could be found in Appendix A.4.
Different platforms have revealed substantial variation in how effectively they address user intentions in multi-turn conversations. ChatGPT and Claude exhibit the highest rates of complete verdicts in Figure 5(a), with both platforms resolving the majority of user intentions fully. Gemini and Grok show more balanced distributions between complete and partial verdicts, with a substantial proportion of intentions receiving incomplete verdicts. Perplexity demonstrates an intermediate pattern with the highest rate of partial verdicts across all platforms. These intention-level patterns are also reflected in the conversation-level completeness scores (Figure 5(b)), where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833.
Also, the number of extracted intentions per conversation (Figure 5(c)) varies systematically across platforms, with ChatGPT and Claude conversations containing a median of 2 intentions, while Gemini, Grok, and Perplexity show a median of 1 intention per conversation, which aligns with the results in Figure 8.
These findings demonstrate the structural diversity of interactions in the dataset, ranging from single-turn queries to complex sessions involving multiple distinct intentions. Furthermore, the prevalence of partial verdicts highlights that many real-world interactions result in incomplete goal satisfaction, a dynamic often obscured in benchmarks that prioritize single-turn correctness over multi-turn resolution. It showcases the functional emphasis of each platform, as the higher intention counts and completeness scores in ChatGPT and Claude reflect their usage for complex or general purpose problem solving, while the frequent partial verdicts and lower intention depth in Perplexity correspond to its search oriented design where model outputs often serve as intermediate information retrieval rather than exhaustive answers. By providing completeness labels alongside the raw conversation data, ShareChat supports the development of automatic evaluation metrics that better align with user satisfaction in multi-turn interactions, which in turn can serve as valuable training signals for reinforcement learning.
4.2 Response Source Analysis
From the statistics, the Grok dataset contains 14,415 conversations, with 8,242 (57.18%) including sources and 6,173 (42.82%) without sources. The Perplexity dataset includes 17,305 conversations, with 8,545 (49.38%) containing sources and 8,760 (50.62%) without sources. More detailed information of these two platforms are provided in Tabel 3.
Plot 6(a) shows that Grok typically uses very few sources per conversation, whereas Perplexity exhibits a long-tailed distribution in which many conversations draw on dozens of sources. In addition, Grok treats “X” as a special source, likely due to privileged access to that platform. This could yield more timely but also potentially noisier or less reliable information, suggesting that the use of X as a primary source warrants further examination for possible misinformation risks. Plot 6(b) further demonstrates Grok’s heavy dependence on this single dominant source (“X”), producing a highly skewed distribution. In contrast, Perplexity draws from a much broader set of sources, with English Wikipedia being the most frequent but not overwhelmingly dominant.
Taken together, these plots demonstrate that the two platforms employ fundamentally distinct information retrieval strategies, prioritizing different types of external evidence to groud their generated responses: Grok gravitates toward X, which is more social-media-driven and volatile, while Perplexity leans toward Wikipedia, which is more curated, structured, and stable.
4.3 Timestamp Analysis
Timestamp data for the Grok and ChatGPT datasets permits a comprehensive temporal analysis as shown in Figure 7. Timestamps were present for the vast majority of turns (99.97% for ChatGPT, 100% for Grok). For the small fraction of missing timestamps in ChatGPT data (1.40% of turns), we employed linear interpolation between known timestamps where feasible, and dropped affected turns when interpolation was not possible. We verified that results remained stable across different handling strategies, confirming that the small amount of missing data did not bias our findings.
After removing extreme outliers above the 99th percentile, we observe systematic differences between platforms. ChatGPT exhibits longer mean user response times than Grok (1,580 vs. 931 seconds), while producing shorter mean LLM response times (18.4 vs. 24.6 seconds). Median values are substantially closer across platforms (users: 93.4 vs. 91.2 seconds; LLMs: 6.9 vs. 16.8 seconds), indicating that mean differences are driven primarily by long-tail behavior.
To assess whether longer LLM outputs demand increased user processing time, we analyze the relationship between LLM response length (in characters) and subsequent user response latency. Across both platforms, correlations are near zero (ChatGPT: ; Grok: ), indicating that verbosity provides little predictive signal for typical user response behavior. While longer responses permit longer delays, they do not systematically increase central tendency, challenging implicit verbosity-as-engagement assumptions. The details could be found in Appendix A.5.
Finally, we examine how response times evolve over the course of a conversation by correlating turn position with response latency. ChatGPT exhibits a significant negative correlation between turn index and LLM response time (Pearson ; Spearman ), indicating progressively faster responses as conversations lengthen, consistent with caching, adaptation, or context optimization effects. In contrast, Grok shows a significant positive correlation (Pearson ; Spearman ), suggesting increasing computational overhead with conversation length. User response times show minimal linear association with turn position on both platforms (ChatGPT: , ; Grok: , ). Together, these divergent temporal dynamics highlight fundamental architectural differences in how the two systems handle extended interactions.
5 Discussion
The findings presented in this study offer several important implications for understanding human-AI interaction and for guiding future research on conversational LLMs. We structure our discussion around four key dimensions enabled by ShareChat. While we only have demonstrated three concrete use cases: conversation completeness, source content analysis and temporal analysis, the dataset’s utility extends well beyond these applications.
Complex Interaction Dynamics in Extended Contexts. ShareChat’s longer conversation contexts, averaging 4.62 turns compared to 2.02 in LMSYS-Chat-1M Zheng et al. (2024), provide a critical resource for studying phenomena that emerge only over extended interactions. Prior work has shown that model reliability can degrade as instructions are refined across turns Laban et al. (2025b), yet most existing corpora offer limited opportunities to investigate such dynamics due to their predominantly single-turn structure. The presence of conversations extending into dozens of turns, combined with the diversity of intention counts per conversation, enables researchers to examine how user needs evolve, how models maintain coherence over long contexts, and when breakdowns occur.
ShareChat makes it possible to study these partial successes systematically, revealing where assistants provide useful but insufficient information and how users adapt their follow-up requests in response. This diversity can inform the development of models better equipped to manage multiple user needs within a single interaction, learning when to shift focus, when to return to unresolved intentions, and how to maintain coherence while serving users whose goals evolve over time.
Platform-Dependent Usage and Performance Variability. Unlike the “universal assistant” assumption often found in general benchmarks, our cross-platform analysis reveals that user behavior is highly sensitive to the specific design and functional positioning of each platform. Our topic distribution analysis demonstrates that users actively select platforms for distinct roles, which aligns their requests with perceived strengths, such as using Perplexity for information retrieval versus Claude for technical reasoning. This functional specialization is mirrored in conversation completeness: while ChatGPT and Claude exhibit narrow, concentrated completeness scores indicating highly predictable output, Gemini and Grok show substantial variability. Notably, the high frequency of “partial” verdicts in Perplexity interactions reflects its distinct role as a search engine, where outputs often serve as intermediate steps rather than final answers. Researchers can use ShareChat to investigate these specialized performance profiles, moving beyond generic comparisons to determine which architectural features drive reliability for specific user intents.
Evidence-Based Reasoning and Temporal Metadata. The preservation of rich metadata enables future research on retrieval-augmented generation (RAG) and interaction pacing in naturalistic settings. Our source content analysis reveals pronounced differences in how platforms integrate external evidence. Grok relies overwhelmingly on concentrated source types, while Perplexity draws from a diversified set of domains. These patterns have implications for understanding how users interpret and rely on cited evidence, and how platform affordances shape trust and verification behaviors. Additionally, the availability of turn-level timestamps enables the investigation of conversational rhythm and interaction pacing, offering a temporal dimension of user experience that text-only corpora cannot provide. Our analysis exposes contrasting architectural dynamics: while ChatGPT demonstrates progressively lower latency as conversations extend over multiple turns, Grok exhibits an inverse trend, with users experiencing slower generation speeds as context length grows.
Self-Selection and Safety in Public Sharing. Finally, the substantially lower toxicity rates in our corpus compared to existing benchmarks, for both user messages and LLM responses, likely reflect the self-selection inherent in public sharing. Users who choose to share conversations may filter for interactions they consider appropriate or valuable, producing a corpus that captures more constructive usage patterns. This suggests that while public datasets may not fully capture the “dark side” of LLM usage (such as jailbreaking attempts), they offer a cleaner, high-quality resource for studying successful task completion and constructive human-AI collaboration.
6 Conclusion
In this study, we addressed the limitations of existing single interface datasets by constructing ShareChat, a large scale, multi platform corpus of user LLM conversations. By compiling more than 140,000 publicly shared interactions from five distinct platforms, we provided a resource that preserves the native metadata and diverse interaction designs of the current AI ecosystem. Our examination of the dataset statistics shows that the self selection process inherent in public sharing leads to a corpus with significantly longer conversation contexts and lower toxicity compared to standard interaction logs.
We further illustrated the value of this data through three specific analysis. First, our conversation completeness analysis indicated that user intentions are more frequently satisfied in general purpose models like ChatGPT and Claude, reflecting their emphasis on broad capabilities. Second, our source content analysis uncovers the distinct information retrieval strategies employed by Grok and Perplexity. Third, our temporal analysis suggests that ChatGPT and Grok exhibit contrasting temporal dynamics. Collectively, these findings highlight pronounced differences in user behavior that align with the functional emphasis of each platform, underscoring the importance of studying LLM capabilities within their specific ecological contexts.
ShareChat offers broader utility for connecting model capabilities with authentic usage patterns. It not only provides a testbed for evaluating model reliability across turns, but also offers rich supervision for training models to handle complex, multi-turn user needs in different settings.
Limitations
Our study is subject to several limitations that should be considered when interpreting the results.
Selection Bias. First, the reliance on publicly shared URLs introduces a trade-off between two forms of bias. On one hand, this approach mitigates the Observer Bias which we talked about in Section 1, where participants alter their behavior simply because they know they are being monitored McCambridge et al. (2014). On the other hand, it introduces Self-Selection Bias Heckman (1979). Because users voluntarily choose which conversations to generate links for, the corpus is likely skewed towards interactions they deem successful, interesting, or novel. Consequently, ShareChat may underrepresent mundane utility tasks or failed interactions where the model performed poorly, as users are less motivated to share such content publicly Hu et al. (2017).
Platform Imbalance. Second, the dataset exhibits significant imbalance across platforms, reflecting current disparities in market share. ChatGPT accounts for over 70% of the collected conversations, whereas Claude represents less than 1%. Although we continue to expand the collection pipeline, researchers should exercise caution when performing direct cross-platform comparisons and may need to employ stratified sampling to account for volume differences.
Proxy Nature of Automated Metrics. Finally, while our completeness analysis utilizes robust LLM-based evaluation, it remains a proxy metric. Automated judges provide a scalable approximation of user satisfaction but cannot perfectly capture the nuanced, subjective quality of human preference. These results should be interpreted as strong signals of structural completeness rather than absolute ground-truth measurements of user happiness. Future research is needed to validate these automated metrics against human judgment and to fully explore the evolving longitudinal dynamics of user behavior.
Preliminary Scope of Analysis. Third, the specific analyses presented in this work, such as the source domain distribution and temporal engagement patterns, are intended as illustrative use cases to demonstrate the dataset’s utility. They are not exhaustive sociological studies. Future research is needed to fully explore these dimensions, particularly how these patterns evolve longitudinally as platform features change.
Ethical Considerations
Data Provenance and Consent. The data collection methodology employed in this study strictly adheres to ethical standards regarding public data usage. The corpus consists exclusively of conversations sourced from URLs that users voluntarily generated and shared via the platforms’ native sharing features. By actively generating a shareable link, users establish a clear intent for public visibility, distinguishing these artifacts from private chat logs or leaked data. We did not attempt to bypass authentication walls, scrape private user accounts, or access data that was not publicly indexed.
Privacy Preservation and De-Identification. Despite the public provenance of the data, we prioritized the protection of user privacy through a rigorous de-identification pipeline designed to mitigate risks associated with inadvertently exposed sensitive information.
PII Redaction: We utilized automated Named Entity Recognition (NER) and regular expression matching to detect and redact Personally Identifiable Information (PII)—including real names, email addresses, phone numbers, and IP addresses—for different languages.
User Identifier Anonymization: We do not store or release original platform-specific user IDs or usernames. All user identifiers were cryptographically hashed and converted into randomized, internal project identifiers during the initial processing stage. This ensures that the published dataset cannot be easily linked back to user profiles on the original platforms.
Aggregated Analysis: The analyses presented in this paper, like the toxicity assessments and completeness scoring, are conducted solely on aggregated statistics. No individual user’s conversation history is singled out for qualitative scrutiny in a manner that could lead to re-identification or personal embarrassment.
Content Safety and Intended Use. Researchers must be aware that ShareChat reflects real-world interactions and, consequently, contains content that may be offensive, biased, or toxic, as quantified in our toxicity analysis. We release this data specifically for research purposes, such as understanding human-AI interaction dynamics, studying safety alignment, and evaluating model performance, and not for training user-facing production models without further filtering. Users of ShareChat are required to adhere to the terms of use, which strictly prohibit attempts to de-anonymize users or leverage the data for malicious surveillance. By limiting collection to publicly accessible links and proactively scrubbing personal data, this work aims to balance the scientific value of ecological data with harm minimization.
Appendix A Appendix
A.1 Turn length Distribution
As shown in Figure 8, both ChatGPT and Claude showing median turn lengths of 2.0, while Gemini, Grok, and Perplexity each maintain a median of 1.0 but with substantially higher variability. The presence of pronounced outliers extending into the hundreds of turns indicates substantial diversity in conversational depth, capturing both brief exchanges and extended dialogues that reflect naturalistic user interactions across platforms.
A.2 Toxicity Analysis Results
Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack. Table 4 presents the mean probability scores for each of these six dimensions by platform. The final column, “Ratio Toxic,” reports the percentage of turns where the general toxicity score exceeded the standard threshold of 0.1.
The results reveal substantial variation across platforms, with Claude exhibiting the highest prevalence of general toxicity with a mean score of 0.0407, followed by Grok at 0.0258, Gemini at 0.0154, ChatGPT at 0.0146, and Perplexity at 0.0088. Table 5 further summarizes these patterns by role, demonstrating that user contributions consistently scored higher than LLM-generated content across the toxicity dimension with 0.0252 versus 0.0168, as well as in specific categories like threat and identity attack. Conversely, LLM responses showed slightly higher mean scores in the obscene dimension, measuring 0.0136 versus 0.0113.
The conversation-level results are shown in Table 6. It was found that toxicity rates vary substantially by detection method, with 6.8% to 9.7% of conversations containing toxic user content and 3.0% to 5.8% containing toxic LLM responses when aggregated across platforms. This higher sensitivity of Detoxify compared to OpenAI is consistent with the patterns observed at the turn level. Furthermore, consistent with the turn-level analysis of platform trends, Perplexity exhibited the lowest prevalence of toxic conversations at roughly 3% to 4%, while Grok generally displayed the highest rates ranging from 11% to 13%. This consistency across both levels of analysis confirms that the observed differences are stable and likely stem from the distinct ways each platform is designed and used.
| Platform | n | Retention % | Toxicity | Severe Tox. | Obscene | Threat | Insult | Identity Attack |
| Claude | 7,455 | 87.7 | 0.0407 | 0.0017 | 0.0297 | 0.0014 | 0.0127 | 0.0009 |
| Grok | 82,689 | 77.9 | 0.0258 | 0.0011 | 0.0115 | 0.0015 | 0.0118 | 0.0022 |
| Gemini | 52,964 | 72.8 | 0.0154 | 0.0010 | 0.0099 | 0.0017 | 0.0098 | 0.0019 |
| ChatGPT | 761,905 | 71.7 | 0.0146 | 0.0006 | 0.0070 | 0.0010 | 0.0075 | 0.0015 |
| Perplexity | 44,474 | 81.8 | 0.0088 | 0.0005 | 0.0040 | 0.0007 | 0.0051 | 0.0014 |
| Metric | N | Retention % | Toxicity | Severe Tox. | Obscene | Threat | Insult | Identity Attack |
| All | 949,487 | 72.9 | 0.0211 | 0.0010 | 0.0124 | 0.0013 | 0.0094 | 0.0016 |
| llm | 471,020 | 72.9 | 0.0168 | 0.0011 | 0.0136 | 0.0006 | 0.0079 | 0.0010 |
| user | 478,467 | 72.9 | 0.0252 | 0.0009 | 0.0113 | 0.0019 | 0.0108 | 0.0021 |
| Role | ChatGPT | Claude | Gemini | Grok | Perplexity | All Platforms | ||||||
| (%) | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI |
| user | 10.7 | 7.5 | 10.2 | 6.1 | 10.1 | 7.9 | 11.0 | 7.1 | 3.7 | 1.7 | 9.7 | 6.8 |
| llm | 3.1 | 6.0 | 4.8 | 5.2 | 2.9 | 6.3 | 5.2 | 8.1 | 0.7 | 2.2 | 3.0 | 5.8 |
| All | 11.6 | 10.2 | 11.1 | 7.7 | 10.6 | 10.3 | 12.9 | 11.0 | 3.9 | 3.0 | 10.6 | 9.4 |
A.3 Detailed Analysis of User Request Topic Distribution
Number of Topics.
Our taxonomy comprises seven high-level topic categories: Multimedia, Other/Unknown, Practical Guidance, Seeking Information, Self-Expression, Technical Help, and Writing. These categories provide a compact yet expressive organization of user intents, enabling consistent comparative analysis across datasets while preserving sufficient semantic granularity for downstream interpretation.
Discovery Prompt.
We provide below the prompt used to discover user-request topics.
Additional Results on Topic Discovery.
We provide detailed statistics across 5 platforms in the Figure 9.
A.4 Conversation Completeness Analysis
We provided the prompts used for the conversation completeness analysis with Qwen3-8B, adapted from the DeepEval framework in this section. To handle conversations that exceed the 40,960-token context limit of Qwen3-8B, we apply token-based truncation to assistant responses while preserving all user content, and we exclude conversations requiring severe truncation (retaining less than 80% of content) to maintain evaluation quality.
Prompt A: User Intention Extraction
Prompt B: Conversation Completeness Labeling
A.5 Additional Temporal Analysis: Text Length and User Response Time

### v2

@@
 Abstract
-While Large Language Models (LLMs) have evolved into distinct platforms with unique interface designs and capabilities, existing public datasets treat models as generic text generators, stripping away the interface context that actively shapes user interaction.
-To address this limitation, we present ShareChat, a large-scale, cross-platform corpus comprising 142,808 conversations and over 660,000 turns collected from publicly shared URLs across five major platforms: ChatGPT, Claude, Gemini, Perplexity, and Grok.
-ShareChat distinguishes itself by preserving native platform affordances often lost in standard logs, including reasoning traces, source links, and code artifacts, while spanning 101 languages over the period from April 2023 to October 2025.
+While academic research typically treats Large Language Models (LLM) as generic text generators, they are distinct commercial products with unique interfaces and capabilities that fundamentally shape user behavior.
+Current datasets obscure this reality by collecting text-only data through uniform interfaces that fail to capture authentic chatbot usage.
+To address this limitation, we present ShareChat, a large-scale corpus of 142,808 conversations (660,293 turns) sourced directly from publicly shared URLs on ChatGPT, Perplexity, Grok, Gemini, and Claude.
+ShareChat distinguishes itself by preserving native platform affordances, such as citations and thinking traces, across a diverse collection covering 101 languages and the period from April 2023 to October 2025.
 Furthermore, ShareChat offers substantially longer context windows and greater interaction depth than prior datasets.
-We demonstrate the dataset’s multifaceted utility through three demonstrative use cases: (1) analyzing conversation completeness to measure user intent satisfaction; (2) evaluating source citation behaviors in content generation; and (3) conducting temporal analysis to track evolving usage patterns.
-This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild
+To illustrate the dataset’s breadth, we present three case studies: a completeness analysis of intent satisfaction, a citation study of model grounding, and a temporal analysis of engagement rhythms.
+This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild.
+The dataset will be publicly available.
 […]
@@
 Yueru Yan Tuc Nguyen Bo Su Melissa Lieffers Thai Le Indiana University Bloomington, USA {yueryan,tucnguye,subo,mealieff,tle}@iu.edu
-1 Introduction
-Conversational large language model (LLM)-based chatbot services have evolved rapidly in the past three years.
-The first widely adopted general purpose LLM chatbot, ChatGPT, was launched in November 2022 and has reached more than 700 million weekly active users by mid 2025 OpenAI Research (2025).
-Following this success, many companies quickly released their own chatbot services, which also attracted substantial public attention.
-For example, Anthropic introduced the Claude families in 2023 Anthropic (2023, 2024, 2025), Google deployed the Gemini family of models Anil et al.
-(2023), xAI launched Grok as a chatbot integrated with the social media platform X xAI (2023), Perplexity emerged as an answer engine that combines conversational interaction with web search Perplexity AI (2024), and DeepSeek released its R1 reasoning model and corresponding consumer chatbot app in January 2025 DeepSeek AI (2025).
-Although all of these services are built on text-based LLMs, they differ in interface design, supported features, and safety policies, which in turn shape how users interact with them.
-For example, Grok can surface live posts from X when providing answers, Claude models are optimized and evaluated for coding, math, and analysis tasks Anthropic (2023, 2024), and Perplexity consistently presents responses with explicit source citations Perplexity AI (2024).
-These platform level design choices are not static.
-Commercial systems typically undergo continuous reinforcement learning and fine tuning on conversation logs and human feedback Ouyang et al.
-(2022); Bai et al.
-(2022); Chen et al.
-(2024), which tends to reinforce platform specific strengths and norms over time.
-As a result, it is important to understand interaction dynamics and user experience across platforms, instead of treating LLM capabilities as a property of models in isolation.
-However, most widely used LLM benchmarks focus on curated single turn tasks, such as multiple choice knowledge tests or coding and math word problems Hendrycks et al.
-(2021); Cobbe et al.
-(2021).
-In practice, the use of conversational LLMs is much more diverse and complex: users ask follow up questions, change goals, or co construct content over many turns.
-Recent work has shown that even state of the art models often get “lost” in multi turn conversations, with reliability degrading as instructions are refined across turns Laban et al.
-(2025a).
-Studying how people actually interact with LLMs in the wild therefore requires large scale, real conversation datasets, rather than only synthetic prompts or short, benchmark style dialogues.
-Several real world chatbot LLM conversation datasets have been released in the past two years Zhao et al.
-(2024); Zheng et al.
-(2024); Köpf et al.
-(2023).
-For instance, WildChat Zhao et al.
-(2024) compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs.
-LMSYS-Chat-1M Zheng et al.
-(2024) contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena).
-OpenAssistant Conversations Köpf et al.
-(2023) provides human generated assistant style dialogues and detailed preference annotations to support alignment research.
-These datasets are important resources, but they have several limitations:
-Single-Interface Mediation: They are all mediated via a central interface even when multiple backend models are available.
-Consequently, they do not fully capture how different commercial platforms, with distinct interaction designs and product positioning, influence the content that users share and the way they engage with LLMs.
-Text-Only Representation: In addition, existing real conversation datasets primarily focus on plain text message content.
-Yet real interactions with contemporary LLM systems can include more structure: users may enable a thinking mode that exposes intermediate reasoning traces, and some platforms attach source links or other metadata to model outputs, such as the cited web pages surfaced by Perplexity or embedded X posts shown by Grok.
-A richer multi platform dataset should preserve these elements when they are available, so that researchers can better understand how users interpret model reasoning, how they rely on external evidence, and how platform affordances shape behavior across turns.
-Limited Interaction Depth: Moreover, although some datasets explicitly include multi turn conversations Zhao et al.
-(2024), the average depth of a conversation is still relatively short.
-Given that context length is central to LLM training and performance, longer real user chatbot conversations are valuable both for understanding how user intents and prompts evolve and for investigating when and how models succeed or fail at using extended context Laban et al.
-(2025a).
-Observer Bias: Finally, data collection in these studies typically relies on pre-interaction consent, where users agree to share their data before or during the interaction in exchange for service access.
-This introduces the Observer Bias (or Hawthorne Effect) McCambridge et al.
-(2014), as users who are aware they are being monitored for research purposes may consciously or unconsciously alter their behavior, which often exhibits higher social desirability compared to authentic, unobserved private usage.
-In contrast, ShareChat relies on post-hoc sharing, where the decision to publish is made only after the natural interaction has concluded, thereby preserving the integrity of the original behavior during the conversation itself.
-In this study, we address these gaps by constructing a large scale, multi platform corpus of authentic user-LLM conversations.
-We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: Claude, Perplexity, Grok, ChatGPT, and Gemini.
-All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than twenty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages.
-The dataset therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers.
-Each conversation record contains the complete sequence of user and assistant turns, along with platform specific metadata such as, when available, timestamps, model thinking traces, or source links attached by the interface.
-The data collection is done under IRB approval (#28569).
-To demonstrate the specific utility of ShareChat, besides describing the basic statistics of the corpus, we present three illustrative case studies, each designed to leverage ShareChat’s unique features to overcome the limitation of prior resources.
-First, we perform conversation completeness analysis, where we label whether specific user intentions are fully addressed.
-This utilizes ShareChat’s high interaction depth to capture the resolution of complex goals often obscured in shorter, shallow exchanges.
-Second, we carry out source domain analysis that connects assistant responses to external resources surfaced by platforms like Perplexity and Grok.
-This leverages our preservation of non-textual metadata to evaluate how models ground their claims, which is a dimension absent in text-only datasets.
-Third, we conduct temporal analysis to examine response time and engagement patterns.
-This exploits our native turn-level timestamps to reveal authentic real-world interaction rhythms, free from the temporal distortions often introduced by observer bias in monitored settings.
-Together, these analyses highlight the value of a multi-platform, multi-turn, and multi-lingual conversation corpus for understanding real user experiences with LLMs.
-Our contributions are twofold.
-First, we construct and analyze what is, to our knowledge, the first multi platform, multi turn, and multi lingual conversation dataset of user LLM interactions, with substantially longer conversations on average than prior corpora and lower levels of toxic content.
-Second, we provide three concrete use cases: conversation completeness analysis, temporal analysis, and source link analysis, which together demonstrate how the dataset can be used to study user interaction behavior in depth.
-2 Data Collection
-For the five differnet platfomrs, we developed systematic, platform specific data extraction pipelines to collect conversation content, metadata, and platform specific features from publicly accessible shared conversation URLs found on the Internet.
-Although we found that such URLs are often shared on social media platforms such as X, Reddit p(formerly Twitter), Reddit, and Discord, exhaustively monitoring these individual ecosystems presents significant scalability challenges.
-To tackle this, we utilized Internet archival services, specifically the Internet Archive (Wayback Machine), to search for and extract URLs matching specific patterns (e.g., “chatgpt.com/share/*” for ChatGPT), This archival-based discovery method is a well-established strategy for aggregating public digital resources, previously employed to compile datasets for web-scale text Raffel et al.
-(2020), historical news archives Hamborg et al.
-(2017), and corporate disclosure analysis Boulland et al.
-(2024).
-Then, for each platform, data collection covered a distinct time window, as summarized in Table 1.
-| Platform | Share URL Format | Collection Period |
-| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
-| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
-| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
-| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
-| Claude | claude.ai/share/* | Not available |
-For every shared conversation page, we used automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation.
-The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform specific metadata.
-When certain elements were only visible after user interaction, we scripted the corresponding actions.
-For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
-Filtering of Identifiable Information.
-We prioritized user privacy by anonymizing personally identifiable information (PII).
-We used Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy222https://spacy.io/ for Named Entity Recognition, and custom rules to identify and remove PII across various data types—such as names, phone numbers, emails, credit cards, and URLs—in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
 | Dataset | #Convs | #Users | #Turns | Avg. Turns | #User Tok | #Chatbot Tok | #Langs |
@@
 | Claude | 946 | – | 4,251 | 4.49 | 19 | ||
+1 Introduction
+Conversational Large Language Model (LLM)-based chatbot services have evolved rapidly in the past three years.
+The first widely adopted general-purpose LLM chatbot, ChatGPT, was launched in November 2022 and has reached more than 700 million weekly active users by mid-2025 openai2025usagepatterns.
+Following this success, many companies quickly released their own chatbot services, which also attracted substantial public attention.
+For example, Anthropic introduced the Claude families in 2023 anthropic2023claude2card; anthropic2024claude3card; Anthropic_2025, Google deployed the Gemini family of models gemini2023report, xAI launched Grok as a chatbot integrated with the social media platform X xai2023grok, and Perplexity emerged as an answer engine that combines conversational interaction with web search perplexityhelp2024.
+Although all of these services are built on text-based LLMs, they differ in interface design, supported features, and safety policies, which in turn shape how users interact with them.
+For example, Grok can surface live posts from X when providing answers, Claude models are optimized and evaluated for coding, math, and analysis tasks anthropic2023claude2card; anthropic2024claude3card, and Perplexity consistently presents responses with explicit source citations perplexityhelp2024.
+Commercial systems typically undergo continuous reinforcement learning and fine tuning on conversation logs and human feedback ouyang2022instructgpt; bai2022hh; chen2024chatgpt, which tends to reinforce platform-specific strengths and norms over time.
+However, current research often fails to capture the complexity of real-world LLM deployment.
+While commercial platforms have evolved into complex ecosystems with unique features, the scientific community relies on datasets that homogenize these interactions.
+As seen in both synthetic xu2024magpie; li2025infinity; ding2023enhancing and real-world datasets zhao2024wildchat; zheng2023lmsys; kopf2023openassistant; zhang2025large; taori2023stanford; conover2023free, current corpora systematically strip away interface context, treating diverse products as identical, generic text boxes.
+For instance, WildChat zhao2024wildchat compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs.
+LMSYS-Chat-1M zheng2023lmsys contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena).
+OpenAssistant Conversations kopf2023openassistant provides human generated assistant style dialogues and detailed preference annotations to support alignment research.
+These datasets are important resources with several limitations:
+1.
+Artificial Uniformity: Existing datasets typically collect data via a single, generic interface to standardize the process.
+This distorts reality.
+By forcing every model into the same neutral interface, these datasets strip away the unique tools that guide user behavior.
+Users do not interact with a generic “model” in the real world; they interact with a specific product, and scientific benchmarks should reflect this distinction.
+2.
+Loss of Non-Textual Context: In addition, while current datasets are largely limited to plain text, authentic LLM interactions rely on rich structural elements, such as visible thinking traces and embedded source citations, that are often stripped from standard logs.
+This prevents researchers from studying how critical product features actually influence user behavior and prompting strategies.
+3.
+Limited Interaction Depth: Moreover, although some datasets explicitly include multi-turn conversations zhao2024wildchat, the average depth of a conversation is still relatively short.
+In practice, the use of conversational LLMs is much more diverse and complex: users ask follow up questions, change goals, or co-construct content over many turns.
+Recent work has shown that even state of the art models often get “lost” in multi-turn conversations, with reliability degrading as instructions are refined across turns laban2025llmslostmultiturnconversation.
+Given that context length is central to LLM training and performance, longer real user-LLM conversations, extended in both turn count and token length, are valuable both for understanding how user intents and prompts evolve and for investigating when and how models succeed or fail laban2025llmslostmultiturnconversation.
+4.
+“Forced” Sharing: Finally, most datasets rely on participants who know they are being watched, introducing the Observer Bias (or Hawthorne Effect) mccambridge2014hawthorne.
+Users who are aware they are being monitored for research purposes may consciously or unconsciously alter their behavior, which often exhibits higher social desirability compared to authentic, unobserved private usage.
+In this study, we address these gaps by constructing a large scale, multi-platform corpus of authentic user-LLM conversations.
+We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: ChatGPT, Perplexity, Grok, Gemini, and Claude, as shown in Tabel 1.
+All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than thirty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages.
+As ShareChat relies on post-hoc sharing, it therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers.
+Each conversation record contains the complete sequence of user and assistant turns, along with platform-specific metadata, when available, timestamps, model thinking traces, or source links attached by the interface.
+Table 2 outlines the distinct data attributes captured across the five platforms, and the details are shown in Appendix A.2.
+The data collection is done under IRB approval (#28569).
+| Feature/Field |
+ChatGPT |
+Perplexity |
+Grok |
+Gemini |
+Claude |
+| Textual Content | ✓ | ✓ | ✓ | ✓ | ✓ |
+| Source Citations | - | ✓ | ✓ | - | - |
+| Thinking Blocks | - | - | ✓ | - | ✓ |
+| Code Artifacts | - | - | - | - | ✓ |
+| Analysis Blocks | - | - | - | - | ✓ |
+| Turn Timestamps | ✓ | - | ✓ | - | - |
+| Model Version | ✓ | - | ✓ | ✓ | - |
+| View/Share Counts | - | ✓ | - | - | - |
+To demonstrate the specific utility of ShareChat, besides describing the basic statistics of the corpus, we present three illustrative case studies.
+First, we perform conversation completeness analysis to assess if user intentions are fully addressed, leveraging ShareChat’s high interaction depth to capture complex goal resolutions.
+Second, we conduct source domain analysis to evaluate how models ground claims, utilizing preserved non-textual metadata to study citation behaviors.
+Third, we execute temporal analysis using turn-level timestamps, to reveal authentic interaction rhythms and engagement patterns free from observer bias.
+Our contributions are twofold.
+First, we construct the first large-scale corpus of user-LLM interactions that preserves native affordances of multiple platforms, with substantially longer conversations on average and lower levels of toxic content than prior corpora.
+Second, we provide three concrete case studies, which together demonstrate how the dataset can be used to study user interaction behavior in depth.
+2 Data Collection
+For the five different platforms, we develop systematic, platform-specific data extraction pipelines to collect conversation content, metadata, and platform-specific features from publicly accessible shared conversation URLs found on the Internet.
+Although we find that such URLs are often shared on social media platforms such as X (formerly Twitter), Reddit, and Discord, exhaustively monitoring these individual ecosystems presents significant scalability challenges.
+To tackle this, we utilize Internet archival services, specifically the Internet Archive (Wayback Machine), to search for and extract URLs matching specific patterns (e.g., “chatgpt.com/share/*” for ChatGPT).
+This archival-based discovery method is a well-established strategy for aggregating public digital resources, previously employed to compile datasets for web-scale text raffel2020t5, historical news archives hamborg2017news, and corporate disclosure analysis boulland2024websites.
+For each platform, data collection covers a distinct time window, as summarized in Appendix A.1.
+For every shared conversation page, we use automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation.
+The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform-specific metadata.
+When certain elements were only visible after user interaction, we script the corresponding actions.
+For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
+Filtering of Identifiable Information.
+We prioritize user privacy by anonymizing personally identifiable information (PII).
+We use Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy honnibal2020spacy for Named Entity Recognition (NER), and custom rules to identify and remove PII across various data types, such as names, phone numbers, emails, credit cards, and URLs, in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
 3 Preliminary Analysis
-The final dataset contains conversations from five platforms, featuring a wide range of languages, diverse user prompts and longer conversation with less toxic content compared to other datasets.
-In this section, we describe the basic statistics of the dataset and platform-specific attributes.
-We also present the results of our toxicity analysis and topic distribution.
+The final dataset contains conversations from five platforms, featuring a wide range of languages, diverse user prompts and longer conversations with less toxic content compared to other datasets.
+In this section, we describe the basic statistics of the dataset, and the results of our toxicity analysis and topic distribution.
 3.1 Basic Statistics
 In total, ShareChat has 142,808 conversations with more than 660,293 interaction turns.
-Table 2 presents comprehensive statistics comparing our multi-platform dataset with existing conversation corpora on number of turns, token length and languages used.
-The language detection was performed at message-level using lingua-py and then combined into conversation-level statistics in Table 2.
-Token counts were computed using the Llama-2 tokenizer for consistent cross-platform comparison Zhao et al.
-(2024).
-In summary, ShareChat significantly exceeds existing corpora in conversation depth, averaging 4.62 turns compared to the 2.02 turns observed in LMSYS-Chat-1M.
-Similarly, it demonstrates greater linguistic diversity by spanning 101 distinct languages, whereas the most diverse competitor covers only 65.
-The model responses are also substantially longer, with a mean of 1115.30 tokens compared to 215.71 for LMSYS-Chat-1M, suggesting the presence of richer and more substantive interactions.
-ChatGPT comprises the largest portion of the corpus with 102,740 conversations, offering the broadest language coverage.
-The key distinguishing characteristic of ShareChat is the substantially longer and more heterogeneous distribution of conversational turns compared to existing public corpora as shown in Figure 1.
-Whereas prominent benchmarks such as Alpaca, Dolly, LMSYS-Chat-1M, and WildChat exhibit median turn lengths of 1.0, our aggregated dataset achieves a median of 2.0 turns, indicating that conversations in our corpus are fundamentally more multi-turn in nature.
-This pattern is consistently maintained across individual platforms with results shown in Appendix A.1.
-For language detection, we first identified the language of each message using the lingua library, and we include these per message labels in the final dataset.
-We then assigned a dominant language to each conversation as the most frequently detected language across the user turns.
-As shown in Figure 2, English accounts for the clear majority of conversations across all platforms, with Japanese as the second most frequent language and all other languages each contributing a much smaller share.
+Table 1 compares ShareChat with existing corpora across key metrics: turn count, token length, and language diversity.
+The language detection was performed at message-level using lingua-py and then combined into conversation-level statistics.
+Token counts were computed using the Llama-2 tokenizer for consistent cross-platform comparison zhao2024wildchat.
+In summary, ShareChat distinguishes itself by extending the scope of analysis along two critical dimensions: interaction depth and content volume.
+First, conversations are substantially longer, averaging 4.62 turns with a heterogeneous distribution that consistently exceeds existing public corpora (Figure 1 and Appendix A.3).
+Second, model responses are significantly denser, with a mean of 1,115.30 tokens compared to just 215.71 for LMSYS-Chat-1M, suggesting the presence of richer and more substantive problem-solving.
+ShareChat also demonstrates greater linguistic diversity by spanning 101 distinct languages, whereas the most diverse competitor covers only 65.
+For languages covered, as shown in Figure 2, English accounts for the clear majority of conversations across all platforms, with Japanese as the second most frequent language and all other languages each contributing a much smaller share.
 Compared to previous datasets, ShareChat has a more balanced language distribution.
-As ShareChat features significantly longer conversation turns and higher token counts per message, it addresses a critical gap in current model development pipelines, which often rely heavily on single-turn instruction tuning.
-By capturing extended interactions, the data supports training for long-context dependencies and exploring complex user behaviors that only emerge during deep, multi-turn exchanges.
-3.2 Platform-Specific Interaction and Metadata Attributes
-Different platforms reflect distinct design philosophies.
-ChatGPT and Gemini offer general-purpose conversational assistance.
-Perplexity focuses on search-oriented information retrieval.
-Grok targets social media-integrated question answering.
-Claude specializes in technical and reasoning support.
-These structural differences align with intended use cases and shape the types of user interactions preserved in ShareChat.
-Table 3 outlines the distinct data attributes captured across the five platforms, which vary significantly based on their functional priorities.
-While ChatGPT and Gemini follow standard conversational formats, other platforms expose specialized content structures.
-Claude is unique in preserving “thinking blocks” and versioned code artifacts, enabling the analysis of intermediate reasoning and iterative coding processes.
-In contrast, Perplexity and Grok structure responses around retrieval.
-Perplexity organizes content into answers, sources, and images with inline citations, while Grok features a dedicated pane for web and social media sources (X posts).
-These structural differences allow for distinct lines of inquiry: technical workflow analysis for Claude versus citation accuracy and source integration studies for Perplexity and Grok.
-The granularity of metadata also dictates the scope of possible temporal analysis.
-ChatGPT and Grok are the only platforms providing turn-level timestamps, which are essential for modeling conversation rhythm and response latency.
-ChatGPT further enriches this with model version identifiers (e.g., gpt-4).
-Conversely, Perplexity lacks turn-level timing but offers unique engagement metrics, including view counts, share counts, and last-updated timestamps, effectively treating conversations as evolving public artifacts rather than static logs.
-| Feature/Field |
-Claude |
-Perplexity |
-Grok |
-ChatGPT |
-Gemini |
-| Basic Turn Structure | ✓ | ✓ | ✓ | ✓ | ✓ |
-| Metadata Capture | ✓ | ✓ | ✓ | ✓ | ✓ |
-| HTML/MHTML | ✓ | ✓ | ✓ | - | ✓ |
-| Code Artifacts | ✓ | - | - | - | - |
-| Thinking Blocks | ✓ | - | ✓ | - | - |
-| Analysis Blocks | ✓ | - | - | - | - |
-| Source Citations | - | ✓ | ✓ | - | - |
-| Image Metadata | - | ✓ | - | - | - |
-| Timestamps / turn | - | - | ✓ | ✓ | - |
-| User Identifiers | ✓ | ✓ | - | ✓ | ✓ |
-| View Counts | - | ✓ | - | - | - |
 3.2 Toxicity Analysis
-To assess toxicity in ShareChat, we employed both DetoxifyHanu and Unitary team (2020) and OpenAI Moderation API333https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat analysis Zhao et al.
-(2024).
+To assess toxicity in ShareChat, we employ both Detoxify hanu2020detoxify and OpenAI Moderation API222https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat zhao2024wildchat.
 Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering.
-Given the language coverage limitations of Detoxify, we retained only conversational turns in languages explicitly supported by the model, resulting in a filtered dataset of 949,487 conversations (72.9% of the original corpus) across all five platforms, with more details shown in Appendix A.2.
-For Detoxify, we applied the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
-The results for turn-level statistics are shown in Figure 3.
-Our analysis revealed substantial differences between the two detection methods, consistent with limited agreement observed in prior work where WildChat Zhao et al.
-(2024) reported that Detoxify flagged 8.12% of user turns compared to the 6.05% flagged by OpenAI Moderation.
-At the turn level, we observed distinct patterns where Detoxify consistently detected higher toxicity rates for user messages, ranging from 2.8% to 5.6%, compared to the 1.3% to 4.0% range identified by OpenAI Moderation across all platforms.
-Notably, Claude showed the highest user toxicity under Detoxify detection at 5.6%.
-There appears to be a positive correlation between user and LLM toxicity rates within platforms.
-For example, Claude exhibited the highest user toxicity with 5.6% via Detoxify and 2.7% via OpenAI, and similarly showed elevated LLM toxicity at 6.4% and 3.1% respectively.
-Conversely, Perplexity demonstrated the lowest rates for both users, at 2.8% via Detoxify and 1.3% via OpenAI, and LLMs, at 0.5% via Detoxify and 1.8% via OpenAI.
-Notably, ShareChat contains substantially less toxic content compared to existing benchmarks.
-Our overall user toxicity rates of 4.1% via Detoxify and 2.9% via OpenAI are considerably lower than both WildChat at 6.05% and LMSYS-Chat-1M at 3.08% using the OpenAI method Zhao et al.
-(2024).
+Given the language coverage limitations of Detoxify, we retain only conversational turns in languages explicitly supported by the model, with more details shown in Appendix A.4.
+For Detoxify, we apply the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
+The results for turn-level toxicity are shown in Figure 3, which reveal substantial differences between the two detection methods, consistent with limited agreement observed in prior work zhao2024wildchat.
+Furthermore, we find a statistically significant positive rank-order correlation between user and LLM toxicity rates within platforms (Spearman’s for Detoxify; for OpenAI).
+This suggests a “mirroring” effect where platforms receiving more toxic user inputs tend to generate more toxic outputs.
+For instance, Claude ranks highest in both user (5.6%) and model toxicity (6.4%) under Detoxify.
+Most importantly, ShareChat contains substantially less toxic content compared to existing benchmarks.
+Our overall user toxicity rates of 4.1% via Detoxify and 2.9% via OpenAI are considerably lower than both WildChat at 6.05% and LMSYS-Chat-1M at 3.08% using the OpenAI method zhao2024wildchat.
 Similarly, our LLM toxicity rates of 1.6% via Detoxify and 3.2% via OpenAI are lower than the 5.18% reported by WildChat and comparable to the 4.12% in LMSYS-Chat-1M.
 3.3 Topic Analysis
-We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 444https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup.
-To enable large-scale processing, we run inference with 4-bit quantization, batch size 32.
+We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 333https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup.
+To enable large-scale processing, we run inference with 4-bit quantization, batch size of 32.
 Model outputs are normalized through a post-processing pipeline to handle formatting variants and stray meta-responses, ensuring reliable mapping to our fixed taxonomy.
 For downstream analyses, the 24 fine-grained categories are further consolidated into seven higher-level groups to support both granular and coarse-grained comparisons across platforms.
-The 7 high-level categories of user requests are plot in the Figure.
-4.
+The 7 high-level categories of user requests are plotted in Figure 4.
 Overall, the averaged distribution reveals that information seeking remains the core driver of LLM interactions, with creative, technical, and advisory uses forming important but secondary dimensions of user behavior.
-Specifically, we observe a strong central tendency in how users engage with large language models.
+Specifically, we observe a strong central tendency in how users engage with LLMs.
 Seeking Information overwhelmingly dominates usage, comprising 39.6% of all requests, underscoring that fact-oriented question-answering remains the primary interaction paradigm.
@@
 Finally, Multimedia accounts for only 2.4% of requests, likely reflecting limited multimodal capabilities at the time of data collection.
-Detailed prompts, category definitions, and classification statistics are provided in the Appendix A.3.
-4 Use Cases
+Also, we find that users employ platforms for distinct roles, such as using Perplexity for “seeking information” versus Claude for “technical thinking”.
+Detailed prompts, category definitions, classification statistics and platform-specific results are provided in the Appendix A.5.
+4 Representative Analyses
 4.1 Conversation Completeness Analysis
-To assess how comprehensively large language models fulfill user needs in multi-turn help-seeking dialogues, we developed an automated completeness analysis pipeline that evaluates whether each user intention across a conversation is adequately addressed by the assistant.
+To assess how comprehensively LLMs fulfill user needs in multi-turn dialogues, we develop an automated completeness analysis pipeline that evaluates whether each user intention across a conversation is adequately addressed by the assistant.
 Our analysis reveals distinct performance across platforms, with ChatGPT and Claude demonstrating the highest consistency in fully satisfying user intentions, while Gemini, Grok and Perplexity exhibit significantly wider variance in conversational quality across multi-turn interactions.
-The pipeline implements a three-stage workflow adapted from the DeepEval555https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework.
-First, we extract user intentions from each conversation by prompting Qwen3-8B Yang et al.
-(2025) with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation.
-Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress).
-Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0.
-This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes.
-We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions.
-The prompts used and experiment details could be found in Appendix A.4.
-Different platforms have revealed substantial variation in how effectively they address user intentions in multi-turn conversations.
-ChatGPT and Claude exhibit the highest rates of complete verdicts in Figure 5(a), with both platforms resolving the majority of user intentions fully.
+To assess the conversation completeness, we implement a three-stage pipeline using Qwen3-8B yang2025qwen3.
+First, we extract distinct user intentions from the conversation history.
+Second, we classify the resolution of each intention as complete, partial, or incomplete.
+Finally, we compute a conversation-level completeness score by weighting these verdicts.
+We sample 100 conversations (20 per platform) for human evaluation.
+Our coding revealed that 85.1% of user intentions are correctly identified, and among these, 92.5% are correctly labeled.
+Through this annotation scheme, the number of extracted intentions per conversation varies systematically across platforms, with ChatGPT and Claude conversations containing a median of 2 intentions, while Gemini, Grok, and Perplexity show a median of 1 intention per conversation.
+Full prompts and implementation details are provided in Appendix A.6.
+Different platforms reveal substantial variation in how effectively they address user intentions in multi-turn conversations.
+ChatGPT and Claude exhibit the highest rates of complete verdicts in Figure 5, with both platforms fully resolving the majority of user intentions.
 Gemini and Grok show more balanced distributions between complete and partial verdicts, with a substantial proportion of intentions receiving incomplete verdicts.
 Perplexity demonstrates an intermediate pattern with the highest rate of partial verdicts across all platforms.
-These intention-level patterns are also reflected in the conversation-level completeness scores (Figure 5(b)), where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833.
-Also, the number of extracted intentions per conversation (Figure 5(c)) varies systematically across platforms, with ChatGPT and Claude conversations containing a median of 2 intentions, while Gemini, Grok, and Perplexity show a median of 1 intention per conversation, which aligns with the results in Figure 8.
-These findings demonstrate the structural diversity of interactions in the dataset, ranging from single-turn queries to complex sessions involving multiple distinct intentions.
-Furthermore, the prevalence of partial verdicts highlights that many real-world interactions result in incomplete goal satisfaction, a dynamic often obscured in benchmarks that prioritize single-turn correctness over multi-turn resolution.
-It showcases the functional emphasis of each platform, as the higher intention counts and completeness scores in ChatGPT and Claude reflect their usage for complex or general purpose problem solving, while the frequent partial verdicts and lower intention depth in Perplexity correspond to its search oriented design where model outputs often serve as intermediate information retrieval rather than exhaustive answers.
-By providing completeness labels alongside the raw conversation data, ShareChat supports the development of automatic evaluation metrics that better align with user satisfaction in multi-turn interactions, which in turn can serve as valuable training signals for reinforcement learning.
-4.2 Response Source Analysis
-From the statistics, the Grok dataset contains 14,415 conversations, with 8,242 (57.18%) including sources and 6,173 (42.82%) without sources.
-The Perplexity dataset includes 17,305 conversations, with 8,545 (49.38%) containing sources and 8,760 (50.62%) without sources.
-More detailed information of these two platforms are provided in Tabel 3.
-Plot 6(a) shows that Grok typically uses very few sources per conversation, whereas Perplexity exhibits a long-tailed distribution in which many conversations draw on dozens of sources.
-In addition, Grok treats “X” as a special source, likely due to privileged access to that platform.
-This could yield more timely but also potentially noisier or less reliable information, suggesting that the use of X as a primary source warrants further examination for possible misinformation risks.
-Plot 6(b) further demonstrates Grok’s heavy dependence on this single dominant source (“X”), producing a highly skewed distribution.
-In contrast, Perplexity draws from a much broader set of sources, with English Wikipedia being the most frequent but not overwhelmingly dominant.
-Taken together, these plots demonstrate that the two platforms employ fundamentally distinct information retrieval strategies, prioritizing different types of external evidence to groud their generated responses: Grok gravitates toward X, which is more social-media-driven and volatile, while Perplexity leans toward Wikipedia, which is more curated, structured, and stable.
+These findings highlight the dataset’s structural diversity, spanning single-turn queries to complex, multi-intention sessions.
+The prevalence of partial verdicts reveals that real-world goals often remain incompletely satisfied, which is a dynamic obscured by benchmarks prioritizing single-turn correctness.
+By providing these completeness labels, ShareChat enables the development of metrics aligned with user satisfaction and offers rich signals for reinforcement learning.
+4.2 Source Grounding Analysis
+We observe distinct information retrieval paradigms between the two search-enabled platforms.
+As detailed in Table 2, external grounding is a core feature for both systems: the Grok dataset contains 14,415 conversations with 8,242 (57.18%) including sources, while the Perplexity dataset includes 17,305 conversations with 8,545 (49.38%) containing sources.
+As illustrated in Figure 6, Grok exhibits a hyper-concentrated reliance on social data.
+Its source distribution is heavily skewed toward the X platform, which accounts for 40,624 references, nearly triple the citations for Wikipedia (12,507).
+This suggests a design optimized for real-time, social-media-driven context, potentially offering higher timeliness but introducing risks related to information stability and bias.
+In contrast, Perplexity demonstrates a diversified, encyclopedic retrieval strategy.
+Its distribution is significantly less concentrated, with English Wikipedia as the leading source (2,919 citations), followed by community forums like Reddit (1,642) and authoritative scientific repositories like NIH (1,339).
+Furthermore, as detailed in Appendix A.7, Perplexity supports much deeper research sessions, with some conversations integrating over 1,000 distinct sources, whereas Grok typically cites fewer than 20 sources per interaction.
+Taken together, these plots demonstrate that the two platforms employ fundamentally distinct information retrieval strategies, prioritizing different types of external evidence to ground their generated responses.
 4.3 Timestamp Analysis
-Timestamp data for the Grok and ChatGPT datasets permits a comprehensive temporal analysis as shown in Figure 7.
-Timestamps were present for the vast majority of turns (99.97% for ChatGPT, 100% for Grok).
-For the small fraction of missing timestamps in ChatGPT data (1.40% of turns), we employed linear interpolation between known timestamps where feasible, and dropped affected turns when interpolation was not possible.
-We verified that results remained stable across different handling strategies, confirming that the small amount of missing data did not bias our findings.
+Timestamp metadata is available for ChatGPT and Grok, covering nearly all turns (99.97% for ChatGPT and 100% for Grok).
+For the small fraction of missing timestamps in ChatGPT data, we employ linear interpolation between known points where feasible, and dropped affected turns when interpolation was not possible.
+We verify that results remained stable across different handling strategies, confirming that the missing data does not bias our findings.
 After removing extreme outliers above the 99th percentile, we observe systematic differences between platforms.
@@
 24.6 seconds).
-Median values are substantially closer across platforms (users: 93.4 vs.
-91.2 seconds; LLMs: 6.9 vs.
-16.8 seconds), indicating that mean differences are driven primarily by long-tail behavior.
-To assess whether longer LLM outputs demand increased user processing time, we analyze the relationship between LLM response length (in characters) and subsequent user response latency.
-Across both platforms, correlations are near zero (ChatGPT: ; Grok: ), indicating that verbosity provides little predictive signal for typical user response behavior.
-While longer responses permit longer delays, they do not systematically increase central tendency, challenging implicit verbosity-as-engagement assumptions.
-The details could be found in Appendix A.5.
-Finally, we examine how response times evolve over the course of a conversation by correlating turn position with response latency.
+Then, we examine how response times evolve over the course of a conversation by correlating turn position with response latency.
 ChatGPT exhibits a significant negative correlation between turn index and LLM response time (Pearson ; Spearman ), indicating progressively faster responses as conversations lengthen, consistent with caching, adaptation, or context optimization effects.
-In contrast, Grok shows a significant positive correlation (Pearson ; Spearman ), suggesting increasing computational overhead with conversation length.
-User response times show minimal linear association with turn position on both platforms (ChatGPT: , ; Grok: , ).
-Together, these divergent temporal dynamics highlight fundamental architectural differences in how the two systems handle extended interactions.
+In contrast, Grok shows a significant positive correlation (Pearson ; Spearman ), suggesting increasing computational overhead as context accumulates.
+User response times show minimal linear association with turn position on both platforms.
+These divergent temporal dynamics highlight fundamental architectural differences in how the two systems handle extended interactions.
+To evaluate whether verbosity alone provides a reliable predictor of user processing time, we additionally examine correlations between LLM message length and subsequent user response latency.
+Figure 7 shows that for both Grok and ChatGPT, mean user response time increases with LLM response length, with the steepest growth at shorter outputs and a plateau emerging beyond approximately 4–6k characters.
+The shaded regions (±1 SD) reveal substantial variability across users.
+Amongst 6k to 10k the variance further widens and this suggests increased heterogeneity in how users engage with lengthy outputs.
+This trend suggests that while longer responses are associated with higher average response times in aggregate, length alone is insufficient to predict individual-level latency, with more details in Appendix A.8.
 5 Discussion
-The findings presented in this study offer several important implications for understanding human-AI interaction and for guiding future research on conversational LLMs.
-We structure our discussion around four key dimensions enabled by ShareChat.
-While we only have demonstrated three concrete use cases: conversation completeness, source content analysis and temporal analysis, the dataset’s utility extends well beyond these applications.
+The findings presented in this study offer several important implications for understanding human-AI interaction and guiding future research on conversational LLMs.
+While we only have demonstrated three concrete representative analyses, the dataset’s utility extends well beyond these applications.
 Complex Interaction Dynamics in Extended Contexts.
-ShareChat’s longer conversation contexts, averaging 4.62 turns compared to 2.02 in LMSYS-Chat-1M Zheng et al.
-(2024), provide a critical resource for studying phenomena that emerge only over extended interactions.
-Prior work has shown that model reliability can degrade as instructions are refined across turns Laban et al.
-(2025b), yet most existing corpora offer limited opportunities to investigate such dynamics due to their predominantly single-turn structure.
+ShareChat’s longer conversation contexts, averaging 4.62 turns compared to 2.02 in LMSYS-Chat-1M zheng2023lmsys, provide a critical resource for studying phenomena that emerge only over extended interactions.
+Prior work has shown that model reliability can degrade as instructions are refined across turns laban2025llms, yet most existing corpora offer limited opportunities to investigate such dynamics due to their predominantly single-turn structure.
 The presence of conversations extending into dozens of turns, combined with the diversity of intention counts per conversation, enables researchers to examine how user needs evolve, how models maintain coherence over long contexts, and when breakdowns occur.
 ShareChat makes it possible to study these partial successes systematically, revealing where assistants provide useful but insufficient information and how users adapt their follow-up requests in response.
-This diversity can inform the development of models better equipped to manage multiple user needs within a single interaction, learning when to shift focus, when to return to unresolved intentions, and how to maintain coherence while serving users whose goals evolve over time.
+This diversity provides the training signals necessary for models to learn adaptive strategies, specifically, when to shift focus, how to revisit unresolved intentions, and how to maintain coherence as user goals evolve.
 Platform-Dependent Usage and Performance Variability.
 Unlike the “universal assistant” assumption often found in general benchmarks, our cross-platform analysis reveals that user behavior is highly sensitive to the specific design and functional positioning of each platform.
-Our topic distribution analysis demonstrates that users actively select platforms for distinct roles, which aligns their requests with perceived strengths, such as using Perplexity for information retrieval versus Claude for technical reasoning.
-This functional specialization is mirrored in conversation completeness: while ChatGPT and Claude exhibit narrow, concentrated completeness scores indicating highly predictable output, Gemini and Grok show substantial variability.
+Our topic distribution analysis demonstrates that users align their requests with perceived platform strengths, such as using Perplexity for information retrieval versus Claude for technical help.
+This functional specialization is mirrored in conversation completeness that while ChatGPT and Claude exhibit narrow, concentrated completeness scores, Gemini and Grok show substantial variability.
 Notably, the high frequency of “partial” verdicts in Perplexity interactions reflects its distinct role as a search engine, where outputs often serve as intermediate steps rather than final answers.
-Researchers can use ShareChat to investigate these specialized performance profiles, moving beyond generic comparisons to determine which architectural features drive reliability for specific user intents.
-Evidence-Based Reasoning and Temporal Metadata.
+Researchers can use ShareChat to investigate these specialized performance profiles, moving beyond generic comparisons to determine which features drive reliability for specific user intents.
+Evidence-Based Generation and Temporal Metadata.
 The preservation of rich metadata enables future research on retrieval-augmented generation (RAG) and interaction pacing in naturalistic settings.
@@
 These patterns have implications for understanding how users interpret and rely on cited evidence, and how platform affordances shape trust and verification behaviors.
-Additionally, the availability of turn-level timestamps enables the investigation of conversational rhythm and interaction pacing, offering a temporal dimension of user experience that text-only corpora cannot provide.
+Additionally, the availability of turn-level timestamps enables the investigation of conversational rhythm and interaction pacing, offering a temporal dimension of user experience.
 Our analysis exposes contrasting architectural dynamics: while ChatGPT demonstrates progressively lower latency as conversations extend over multiple turns, Grok exhibits an inverse trend, with users experiencing slower generation speeds as context length grows.
@@
 Users who choose to share conversations may filter for interactions they consider appropriate or valuable, producing a corpus that captures more constructive usage patterns.
-This suggests that while public datasets may not fully capture the “dark side” of LLM usage (such as jailbreaking attempts), they offer a cleaner, high-quality resource for studying successful task completion and constructive human-AI collaboration.
+This suggests that ShareChat offers a cleaner, high-quality resource for studying successful task completion and constructive human-AI collaboration.
 6 Conclusion
-In this study, we addressed the limitations of existing single interface datasets by constructing ShareChat, a large scale, multi platform corpus of user LLM conversations.
-By compiling more than 140,000 publicly shared interactions from five distinct platforms, we provided a resource that preserves the native metadata and diverse interaction designs of the current AI ecosystem.
-Our examination of the dataset statistics shows that the self selection process inherent in public sharing leads to a corpus with significantly longer conversation contexts and lower toxicity compared to standard interaction logs.
-We further illustrated the value of this data through three specific analysis.
-First, our conversation completeness analysis indicated that user intentions are more frequently satisfied in general purpose models like ChatGPT and Claude, reflecting their emphasis on broad capabilities.
+In this study, we address the limitations of existing datasets by constructing ShareChat, a large-scale, multi-platform corpus of user-LLM conversations.
+By compiling more than 140,000 publicly shared interactions from five distinct platforms, we provide a resource that preserves the native metadata and diverse interaction designs of the current AI ecosystem.
+The dataset’s basic statistics show that the self-selection process inherent in public sharing leads to a corpus with significantly longer conversation contexts and lower toxicity.
+We further illustrate the value of this data through three specific analyses.
+First, our conversation completeness analysis indicate that user intentions are more frequently satisfied in general purpose models like ChatGPT and Claude, reflecting their emphasis on broad capabilities.
 Second, our source content analysis uncovers the distinct information retrieval strategies employed by Grok and Perplexity.
 Third, our temporal analysis suggests that ChatGPT and Grok exhibit contrasting temporal dynamics.
-Collectively, these findings highlight pronounced differences in user behavior that align with the functional emphasis of each platform, underscoring the importance of studying LLM capabilities within their specific ecological contexts.
-ShareChat offers broader utility for connecting model capabilities with authentic usage patterns.
-It not only provides a testbed for evaluating model reliability across turns, but also offers rich supervision for training models to handle complex, multi-turn user needs in different settings.
+Collectively, these findings highlight pronounced differences in user behavior that align with the functional emphasis of each platform, underscoring the importance of studying LLM capabilities within their specific contexts.
+ShareChat not only provides a testbed for evaluating model reliability across turns, but also offers rich supervision for training models to handle complex, multi-turn user needs in different settings.
 Limitations
@@
 First, the reliance on publicly shared URLs introduces a trade-off between two forms of bias.
-On one hand, this approach mitigates the Observer Bias which we talked about in Section 1, where participants alter their behavior simply because they know they are being monitored McCambridge et al.
-(2014).
-On the other hand, it introduces Self-Selection Bias Heckman (1979).
-Because users voluntarily choose which conversations to generate links for, the corpus is likely skewed towards interactions they deem successful, interesting, or novel.
-Consequently, ShareChat may underrepresent mundane utility tasks or failed interactions where the model performed poorly, as users are less motivated to share such content publicly Hu et al.
-(2017).
+On one hand, this approach mitigates the Observer Bias which we talked about in Section 1.
+On the other hand, it introduces Self-Selection Bias heckman1979sample, because users voluntarily choose which conversations to generate links for, the corpus is likely skewed towards interactions they deem successful, interesting, or novel.
+Therefore, ShareChat may underrepresent mundane utility tasks or failed interactions where the model performed poorly, as users are less motivated to share such content publicly hu2017selection.
 Platform Imbalance.
@@
 Proxy Nature of Automated Metrics.
-Finally, while our completeness analysis utilizes robust LLM-based evaluation, it remains a proxy metric.
+Third, while our conversation completeness analysis and topic analysis utilize robust LLM-based evaluation, it remains a proxy metric.
 Automated judges provide a scalable approximation of user satisfaction but cannot perfectly capture the nuanced, subjective quality of human preference.
@@
 Preliminary Scope of Analysis.
-Third, the specific analyses presented in this work, such as the source domain distribution and temporal engagement patterns, are intended as illustrative use cases to demonstrate the dataset’s utility.
+Finally, the specific analyses presented in this work, such as the source domain distribution and temporal engagement patterns, are intended as illustrative representative analyses to demonstrate the dataset’s utility.
 They are not exhaustive sociological studies.
@@
 Despite the public provenance of the data, we prioritized the protection of user privacy through a rigorous de-identification pipeline designed to mitigate risks associated with inadvertently exposed sensitive information.
-PII Redaction: We utilized automated Named Entity Recognition (NER) and regular expression matching to detect and redact Personally Identifiable Information (PII)—including real names, email addresses, phone numbers, and IP addresses—for different languages.
-User Identifier Anonymization: We do not store or release original platform-specific user IDs or usernames.
-All user identifiers were cryptographically hashed and converted into randomized, internal project identifiers during the initial processing stage.
-This ensures that the published dataset cannot be easily linked back to user profiles on the original platforms.
-Aggregated Analysis: The analyses presented in this paper, like the toxicity assessments and completeness scoring, are conducted solely on aggregated statistics.
+First, we utilized automated NER and regular expression matching to detect and redact PII, including real names, email addresses, phone numbers, and IP addresses, for different languages.
+In addition, we do not store or release original platform-specific user IDs or usernames.
+All user identifiers were cryptographically hashed and converted into internal project identifiers during the initial processing stage.
+Finally, the analyses presented in this paper, like the toxicity assessments and completeness scoring, are conducted solely on aggregated statistics.
 No individual user’s conversation history is singled out for qualitative scrutiny in a manner that could lead to re-identification or personal embarrassment.
@@
 Appendix A Appendix
-A.1 Turn length Distribution
+A.1 Data collection
+Table 3 details the target URLs used for extraction and the data collection timeframe for each platform.
+| Platform | Share URL Format | Collection Period |
+| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
+| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
+| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
+| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
+| Claude | claude.ai/share/* | Not available |
+A.2 Platform-Specific Interaction and Metadata Attributes
+Different platforms also reflect distinct design philosophies, ChatGPT and Gemini offer general-purpose conversational assistance.
+Perplexity focuses on search-oriented information retrieval.
+Grok targets social media-integrated question answering.
+Claude specializes in technical and reasoning support.
+These structural differences align with intended representative analyses and shape the types of user interactions preserved in ShareChat.
+As shown in Table 2, while ChatGPT and Gemini follow standard conversational formats, other platforms expose specialized content structures.
+Claude is unique in preserving “thinking blocks” and versioned code artifacts, enabling the analysis of intermediate reasoning and iterative coding processes.
+In contrast, Perplexity and Grok structure responses around retrieval.
+Perplexity organizes content into answers, sources, and images with inline citations, while Grok features a dedicated pane for web and social media sources (X posts).
+These structural differences allow for distinct lines of inquiry: technical workflow analysis for Claude versus citation accuracy and source integration studies for Perplexity and Grok.
+The granularity of metadata also dictates the scope of possible temporal analysis.
+ChatGPT and Grok are the only platforms providing turn-level timestamps, which are essential for modeling conversation rhythm and response latency.
+ChatGPT further enriches this with model version identifiers (e.g., gpt-4).
+Conversely, Perplexity lacks turn-level timing but offers unique engagement metrics, including view counts, share counts, and last-updated timestamps, effectively treating conversations as evolving public artifacts rather than static logs.
+A.3 Turn length Distribution
 As shown in Figure 8, both ChatGPT and Claude showing median turn lengths of 2.0, while Gemini, Grok, and Perplexity each maintain a median of 1.0 but with substantially higher variability.
 The presence of pronounced outliers extending into the hundreds of turns indicates substantial diversity in conversational depth, capturing both brief exchanges and extended dialogues that reflect naturalistic user interactions across platforms.
-A.2 Toxicity Analysis Results
+A.4 Toxicity Analysis Results
 Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack.
+After filtering conversations supported by the model, resulting in a filtered dataset of 104,107 (72.9% of the original corpus) across all five platforms.
 Table 4 presents the mean probability scores for each of these six dimensions by platform.
@@
 | Role | ChatGPT | Claude | Gemini | Grok | Perplexity | All Platforms | ||||||
-| (%) | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI | Detox | OpenAI |
+| (%) | Detox | Open | Detox | Open | Detox | Open | Detox | Open | Detox | Open | Detox | Open |
 | user | 10.7 | 7.5 | 10.2 | 6.1 | 10.1 | 7.9 | 11.0 | 7.1 | 3.7 | 1.7 | 9.7 | 6.8 |
@@
 | All | 11.6 | 10.2 | 11.1 | 7.7 | 10.6 | 10.3 | 12.9 | 11.0 | 3.9 | 3.0 | 10.6 | 9.4 |
-A.3 Detailed Analysis of User Request Topic Distribution
+A.5 Detailed Analysis of User Request Topic Distribution
 Number of Topics.
@@
 We provide detailed statistics across 5 platforms in the Figure 9.
-A.4 Conversation Completeness Analysis
-We provided the prompts used for the conversation completeness analysis with Qwen3-8B, adapted from the DeepEval framework in this section.
+A.6 Conversation Completeness Analysis
+The pipeline implements a three-stage workflow adapted from the DeepEval444https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework.
+First, we extract user intentions from each conversation by prompting Qwen3-8B yang2025qwen3 with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation.
+Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress).
+Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0.
+This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes.
+We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions.
+During manual coding, we observe that mis-classification often occurred when users pasted text, such as articles or messages, without explicit instructions or questions.
+In these instances, the LLM generate responses based solely on an inferred understanding of the content.
+Figure 10 shows the conversation-level completeness scores, where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833.
+Figure 11 shows the number of interactions across platforms.
+The prompts used are as follows.
 To handle conversations that exceed the 40,960-token context limit of Qwen3-8B, we apply token-based truncation to assistant responses while preserving all user content, and we exclude conversations requiring severe truncation (retaining less than 80% of content) to maintain evaluation quality.
@@
 Prompt B: Conversation Completeness Labeling
-A.5 Additional Temporal Analysis: Text Length and User Response Time
+A.7 Additional Source Grounding Analysis
+To understand the retrieval intensity of search-enabled platforms, we analyze the distribution of source counts per conversation.
+Figure 12 shows that Grok typically uses very few sources per conversation, whereas Perplexity exhibits a long-tailed distribution in which many conversations draw on dozens of sources.
+A.8 Additional Temporal Analysis
+We expand on the temporal dynamics of the dataset by examining both the diurnal rhythms of user activity and the granular relationship between content length and reading time.
+First, to understand when users are most active, Figure 13 presents the normalized hourly activity for each platform, with timestamps adjusted to the users’ local time.
+Second, we investigate whether longer model outputs consistently drive longer user engagement.
+Figure 14 plots the raw distribution of user response times against LLM response length.
+The scatter plot reveals a high degree of variance, with Pearson correlations near zero for both platforms (ChatGPT: ; Grok: ).
+This indicates that while aggregate trends exist, model verbosity is a poor predictor of dwell time.

### v3

@@
 This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild.
-The dataset will be publicly available.
+The dataset is publicly available via Hugging Face 111https://github.com/raye22/ShareChat222https://huggingface.co/datasets/tucnguyen/ShareChat.
 […]
@@
 | LMSYS-Chat-1M | 1,000,000 | 210,479 | 2,020,000 | 2.02 | 65 | ||
+| WILDCHAT | 1,039,785 | 204,736 | 2,641,054 | 2.54 | 68 | ||
 | Our Multi-Platform Dataset | |||||||
@@
 While commercial platforms have evolved into complex ecosystems with unique features, the scientific community relies on datasets that homogenize these interactions.
-As seen in both synthetic xu2024magpie; li2025infinity; ding2023enhancing and real-world datasets zhao2024wildchat; zheng2023lmsys; kopf2023openassistant; zhang2025large; taori2023stanford; conover2023free, current corpora systematically strip away interface context, treating diverse products as identical, generic text boxes.
+As seen in both synthetic xu2024magpie; li2025infinity; ding2023enhancing and real-world datasets including WildChat zhao2024wildchat, LMSYS-Chat-1M zheng2023lmsys, OpenAssistant kopf2023openassistant, Alpaca taori2023stanford, ShareGPT333https://sharegpt.com/, and Dolly conover2023free, current corpora strip away interface context, treating diverse products as identical, generic text boxes.
 For instance, WildChat zhao2024wildchat compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs.
@@
 Table 2 outlines the distinct data attributes captured across the five platforms, and the details are shown in Appendix A.2.
-The data collection is done under IRB approval (#28569).
+Data collection is done with IRB approval (#28569).
 | Feature/Field |
@@
 For each platform, data collection covers a distinct time window, as summarized in Appendix A.1.
-For every shared conversation page, we use automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation.
+For every shared conversation page, we use an automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation.
 The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform-specific metadata.
@@
 We prioritize user privacy by anonymizing personally identifiable information (PII).
-We use Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy honnibal2020spacy for Named Entity Recognition (NER), and custom rules to identify and remove PII across various data types, such as names, phone numbers, emails, credit cards, and URLs, in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
+We use Microsoft’s Presidio444https://microsoft.github.io/presidio/ as the framework to remove PII across various data types, such as names, phone numbers, emails, credit cards, driver licenses and URLs, in multiple languages, including English, Spanish, German, French, Italian, Portuguese, Dutch, Chinese, Japanese, Russian, and Hebrew.
+To ensure rigorous privacy preservation, we employed an automated LLM-based evaluator to audit the PII removal process, with detailed results provided in Appendix A.9.
 3 Preliminary Analysis
@@
 3.2 Toxicity Analysis
-To assess toxicity in ShareChat, we employ both Detoxify hanu2020detoxify and OpenAI Moderation API222https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat zhao2024wildchat.
+To assess toxicity in ShareChat, we employ both Detoxify hanu2020detoxify and OpenAI Moderation API555https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat zhao2024wildchat.
 Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering.
@@
 3.3 Topic Analysis
-We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 333https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup.
+We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 666https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup.
 To enable large-scale processing, we run inference with 4-bit quantization, batch size of 32.
@@
 As detailed in Table 2, external grounding is a core feature for both systems: the Grok dataset contains 14,415 conversations with 8,242 (57.18%) including sources, while the Perplexity dataset includes 17,305 conversations with 8,545 (49.38%) containing sources.
-As illustrated in Figure 6, Grok exhibits a hyper-concentrated reliance on social data.
+As shown in Figure 6, the source of the name is extracted from the website’s top-level domain.
+Only top 10 of each platform are displayed in the rose graph.
+Grok exhibits a hyper-concentrated reliance on social data.
 Its source distribution is heavily skewed toward the X platform, which accounts for 40,624 references, nearly triple the citations for Wikipedia (12,507).
@@
 Table 3 details the target URLs used for extraction and the data collection timeframe for each platform.
-| Platform | Share URL Format | Collection Period |
-| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
-| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
-| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
-| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
-| Claude | claude.ai/share/* | Not available |
+Creation timestamps were not captured for the Claude platform in this iteration; however, we plan to incorporate this metadata in future versions of the dataset.
 A.2 Platform-Specific Interaction and Metadata Attributes
@@
 The presence of pronounced outliers extending into the hundreds of turns indicates substantial diversity in conversational depth, capturing both brief exchanges and extended dialogues that reflect naturalistic user interactions across platforms.
-A.4 Toxicity Analysis Results
-Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack.
-After filtering conversations supported by the model, resulting in a filtered dataset of 104,107 (72.9% of the original corpus) across all five platforms.
-Table 4 presents the mean probability scores for each of these six dimensions by platform.
-The final column, “Ratio Toxic,” reports the percentage of turns where the general toxicity score exceeded the standard threshold of 0.1.
-The results reveal substantial variation across platforms, with Claude exhibiting the highest prevalence of general toxicity with a mean score of 0.0407, followed by Grok at 0.0258, Gemini at 0.0154, ChatGPT at 0.0146, and Perplexity at 0.0088.
-Table 5 further summarizes these patterns by role, demonstrating that user contributions consistently scored higher than LLM-generated content across the toxicity dimension with 0.0252 versus 0.0168, as well as in specific categories like threat and identity attack.
-Conversely, LLM responses showed slightly higher mean scores in the obscene dimension, measuring 0.0136 versus 0.0113.
-The conversation-level results are shown in Table 6.
-It was found that toxicity rates vary substantially by detection method, with 6.8% to 9.7% of conversations containing toxic user content and 3.0% to 5.8% containing toxic LLM responses when aggregated across platforms.
-This higher sensitivity of Detoxify compared to OpenAI is consistent with the patterns observed at the turn level.
-Furthermore, consistent with the turn-level analysis of platform trends, Perplexity exhibited the lowest prevalence of toxic conversations at roughly 3% to 4%, while Grok generally displayed the highest rates ranging from 11% to 13%.
-This consistency across both levels of analysis confirms that the observed differences are stable and likely stem from the distinct ways each platform is designed and used.
 | Platform | n | Retention % | Toxicity | Severe Tox. | Obscene | Threat | Insult | Identity Attack |
@@
 | All | 11.6 | 10.2 | 11.1 | 7.7 | 10.6 | 10.3 | 12.9 | 11.0 | 3.9 | 3.0 | 10.6 | 9.4 |
+A.4 Toxicity Analysis Results
+Leveraging the multi-label capabilities of the Detoxify model trained on the Jigsaw Toxic Comment Classification dataset, we analyzed six distinct dimensions of harmful content: toxicity, which is a broad category for rude or disrespectful comments, severe toxicity, obscenity, threat, insult, and identity attack.
+After filtering conversations supported by the model, resulting in a filtered dataset of 104,107 (72.9% of the original corpus) across all five platforms.
+Table 4 presents the mean probability scores for each of these six dimensions by platform.
+The results reveal substantial variation across platforms, with Claude exhibiting the highest prevalence of general toxicity with a mean score of 0.0407, followed by Grok at 0.0258, Gemini at 0.0154, ChatGPT at 0.0146, and Perplexity at 0.0088.
+Table 5 further summarizes these patterns by role, demonstrating that user contributions consistently scored higher than LLM-generated content across the toxicity dimension with 0.0252 versus 0.0168, as well as in specific categories like threat and identity attack.
+Conversely, LLM responses showed slightly higher mean scores in the obscene dimension, measuring 0.0136 versus 0.0113.
+The conversation-level results are shown in Table 6.
+It was found that toxicity rates vary substantially by detection method, with 6.8% to 9.7% of conversations containing toxic user content and 3.0% to 5.8% containing toxic LLM responses when aggregated across platforms.
+This higher sensitivity of Detoxify compared to OpenAI is consistent with the patterns observed at the turn level.
+Furthermore, consistent with the turn-level analysis of platform trends, Perplexity exhibited the lowest prevalence of toxic conversations at roughly 3% to 4%, while Grok generally displayed the highest rates ranging from 11% to 13%.
+This consistency across both levels of analysis confirms that the observed differences are stable and likely stem from the distinct ways each platform is designed and used.
 A.5 Detailed Analysis of User Request Topic Distribution
@@
 A.6 Conversation Completeness Analysis
-The pipeline implements a three-stage workflow adapted from the DeepEval444https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework.
+The pipeline implements a three-stage workflow adapted from the DeepEval777https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework.
 First, we extract user intentions from each conversation by prompting Qwen3-8B yang2025qwen3 with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation.
@@
 In these instances, the LLM generate responses based solely on an inferred understanding of the content.
-Figure 10 shows the conversation-level completeness scores, where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833.
+Figure 10 shows the conversation-level completeness scores, where ChatGPT, Claude, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833 and Grok’s median is 0.900.
 Figure 11 shows the number of interactions across platforms.
@@
 This indicates that while aggregate trends exist, model verbosity is a poor predictor of dwell time.
+A.9 PII Removal Validation and Data Filtering
+To validate the efficacy of our de-identification pipeline, we employed GPT-OSS-120B, accessed through the Jetstream Cloud API888https://docs.jetstream-cloud.org/inference-service/overview/, as an automated auditor to detect any residual personally identifiable information in the anonymized text.
+The model was configured with medium reasoning effort to optimize for computational efficiency while maintaining evaluation quality.
+The quantitative results of this automated audit are presented in Table 7.
+To further verify these automated findings, we conducted a manual review of 50 randomly sampled entries (228 turns) that the evaluator flagged as containing PII (i.e., "Positive" labels).
+This human verification revealed a true positive rate of only 3.7%, indicating that the vast majority of flagged instances were false positives and that the pipeline is highly effective at removing sensitive data.
+Consequently, to ensure the highest privacy standards, the final released dataset excludes languages not explicitly supported by the Presidio framework.
+However, for conversations in these unsupported languages, we provide the original source URLs to facilitate future research while maintaining the privacy integrity of the main corpus.
+The final composition of the filtered dataset, utilizing strict language and PII filtering, is detailed in Table 8.
+| Platform | Success Rate | Flagged | Total |
+| ChatGPT | 95.20% | 51041 | 1062949 |
+| Perplexity | 94.42% | 2,899 | 54,355 |
+| Grok | 94.15% | 6,010 | 106,168 |
+| Gemini | 95.43% | 3,302 | 72,746 |
+| Claude | 97.01% | 252 | 8,504 |
+| Platform | Original | Removed | % Kept |
+| ChatGPT | 1,062,949 | 86,571 | 91.9% |
+| Perplexity | 54,355 | 5,754 | 89.4% |
+| Grok | 106,168 | 7,280 | 93.1% |
+| Gemini | 72,746 | 5,078 | 93.0% |
+| Claude | 8,504 | 140 | 98.4% |
+| Total | 1,304,722 | 104,823 | 92.0% |